# Improved Trade-offs Between Amortization and Download Bandwidth for Linear HSS

### v1

@@ 4.1 Hermitian Code Definition, Parameters
:=$\mathcal{H}:=\left\{\mathrm{ev}_{\mathcal{Z}}(f)\;:\;f\in M\right\}.$
@@
The $k$-dimensional Hermitian code $\mathcal{H}$ is a linear code of length $n=q^{3}$ and minimum distance
@@ 5 Linear HSS from Goppa Codes
Then there exists a $t$-private, $s$-server HSS $\pi=\left(\mathsf{Share},\mathsf{Eval},\mathsf{Rec}\right)$ for some non-trivial $\mathcal{F}\subseteq\mathrm{POLY}_{d,m}(\mathbb{F}_{2})$ with
@@
Figure 2: The left (right) plot compares the download rates (amortization parameters) of [FIKW22], [BW23] with that achieved by Theorem 19 when $d=t=2$. The $x$-axis represents the number of servers and ranges from 1 to 512. This emphasizes the order-of-magnitude amortization savings of Theorem 19 at practical parameter regimes relative to [FIKW22], [BW23], with small concessions to rate.
@@ 5.1 Goppa Code Definition, Parameters
For some $n\in\mathbb{Z}^{+}$, fix $V=\{\alpha_{1},\ldots,\alpha_{n}\}\subseteq\mathbb{F}_{2^{u}},\;u\in\mathbb{Z}^{+}$. The Goppa polynomial is a polynomial
$g(x)=g_{V}(s)=g_{0}+g_{1}x+\cdots+g_{r}x^{r}\in\mathbb{F}_{q^{u}}[x]$
@@
In practice, it is common to take $n=2^{u}$ and set $V=\mathbb{F}_{2^{u}}$ for some $u\in\mathbb{Z}^{+}$, as it reduces the task of finding a suitable Goppa polynomial to that of finding a irreducible polynomial of degree $r$ over $\mathbb{F}_{2^{u}}$.
@@
For $n,u,r\in\mathbb{Z}^{+}$ let $\Gamma=\Gamma_{n,u,r}$ be a Goppa code as in Definition 11. Then $\Gamma$ is a linear code of length $n$, dimension $k\geq n-ur$, and minimum distance $d(\Gamma)\geq r+1$.

### v2

@@ 4.1 Hermitian Code Definition, Parameters
:=$\mathcal{H}:=\left\{\mathrm{ev}_{\mathcal{Z}}(f)\;:\;f\in M\right\}$
where $\mathrm{ev}_{\mathcal{Z}}(f)=\left(f(x,y):(x,y)\in\mathcal{Z}\right)$ denotes the standard evaluation map.
@@
The $k$-dimensional Hermitian code $\mathcal{H}$ is a linear code of length $n=q^{3}$ and rate $k/n$ with minimum distance
@@ 5 Linear HSS from Goppa Codes
Then there exists an explicit $t$-private, $s$-server HSS $\pi=\left(\mathsf{Share},\mathsf{Eval},\mathsf{Rec}\right)$ for some non-trivial $\mathcal{F}\subseteq\mathrm{POLY}_{d,m}(\mathbb{F}_{2})$ with
@@
Figure 2: The left (right) plot compares the download rates (amortization parameters) of [FIKW22], [BW23] with that achieved by Theorem 19 when $d=t=2$. The $x$-axis represents the number of servers and ranges from 1 to 512. This emphasizes the super-constant amortization savings of Theorem 19 at practical parameter regimes relative to [FIKW22], [BW23], with small concessions to rate.
@@ 5.1 Goppa Code Definition, Parameters
For some $n\in\mathbb{Z}^{+}$, fix $V=\{\alpha_{1},\ldots,\alpha_{n}\}\subseteq\mathbb{F}_{2^{u}},\;u\in\mathbb{Z}^{+}$. A Goppa polynomial is a polynomial
$g(x)=g_{V}(x)=g_{0}+g_{1}x+\cdots+g_{r}x^{r}\in\mathbb{F}_{2^{u}}[x]$
@@
In practice, it is common to take $n=2^{u}$ and set $V=\mathbb{F}_{2^{u}}$ for some $u\in\mathbb{Z}^{+}$, as it reduces the task of finding a suitable Goppa polynomial to that of finding an irreducible polynomial of degree $r$ over $\mathbb{F}_{2^{u}}$.
@@
For $n,u,r\in\mathbb{Z}^{+}$ let $\Gamma=\Gamma_{n,u,r}$ be a binary Goppa code as in Definition 11. Then $\Gamma$ is a linear code of length $n$, dimension $k\geq n-ur$, and minimum distance $d(\Gamma)\geq r+1$.